OpenStudy (anonymous):

g(t)=√3:(t-63) answer should be in interval notation

OpenStudy (anonymous):

is it\[\sqrt{\frac{3}{t-63}}\]?

OpenStudy (anonymous):

yeah.

OpenStudy (anonymous):

the 3 is small above the square root symbol

OpenStudy (amistre64):

3rt(t-63) ; or cbrt(t-63) are pretty good keyboard notation for indexes

OpenStudy (anonymous):

so its\[\sqrt[3]{t-63}\]

OpenStudy (amistre64):

what defines an answer in order to put it into interval notation ???

OpenStudy (anonymous):

it can't be zero underneath the sqrt right?

OpenStudy (amistre64):

odd functions are allowed to go negative

OpenStudy (amistre64):

cbrt(-8) = -2 since, -2*-2*-2 = (-2)^3 = -8

OpenStudy (anonymous):

and this is an odd function?

OpenStudy (amistre64):

is 3 and odd number?

OpenStudy (anonymous):

haha yeah.

OpenStudy (amistre64):

then yes :)

OpenStudy (amistre64):

you looking for domain, or range?

OpenStudy (anonymous):

domain :)

OpenStudy (amistre64):

and we aint got any fractions do we ?

OpenStudy (amistre64):

its just t-63 under there; and not 1/(t-63) right?

OpenStudy (anonymous):

yeah just t-63

OpenStudy (amistre64):

then there are no restrictions to be had; all real numbers can be used domain is then \((-\infty,\infty)\)

OpenStudy (anonymous):

how do you know for sure whether there are restrictions or not? I have a hard time determining this...

OpenStudy (amistre64):

there are a few things to keep in mind; you cant divide by zero, you cant have logs that are less than or equal to zero .... even radicals cant be negative.

OpenStudy (amistre64):

thats pretty much it ....

OpenStudy (anonymous):

okay gotcha. so, would f(x)=sqrt(73.5-7x)... I solved and somehow got 10.5..?

OpenStudy (amistre64):

is sqrt and even or odd radical?

OpenStudy (anonymous):

even..?

OpenStudy (amistre64):

correct , 2 is even :) so, we cant have negative values; so determine the solution to:\[73.5-7x >= 0\]

OpenStudy (anonymous):

And I got 

OpenStudy (amistre64):

nice

OpenStudy (amistre64):

so, lets use 10.5, since we can be equal to it we use the = notation connected by a bar